OpenStudy (anonymous):

Suppose a pick-5 lottery is played as follows. Each ticket (panel) has two sets of numbers.  The top half of the panel has numbers 1 through 40. You must choose 4 distinct numbers out of these, called the White Balls.  The bottome half has numbers 1 through 20. You must choose 1 number out of these, called the Power-Ball. Compute each of the following. Show the combinatorial derivation. Show each probabilty in the form 1=x where x is the nearest integer.

OpenStudy (anonymous):

(a) Total number of distinct ways to play a ticket. (b) Probability of winning the Jackpot. (Match all 4 White Balls, and the Power Ball.) (c) Probability of matching 3 of the White Ball, and the Power Ball. (d) Probability of matching 3 of the White Balls only any one quick please hv exam in few mins!!

OpenStudy (anonymous):

@jim_thompson5910 and @phi any ideas guys ?

jimthompson5910 (jim_thompson5910):

hmm sorry, I'd have to give it more thought and I wish I had more time (instead of a few mins)

jimthompson5910 (jim_thompson5910):

for a) it sounds like it would be (40 C 4)*20 = 91390*20 = 1,827,800 but not 100% sure

OpenStudy (anonymous):

hv 30 mins :(